OpenStudy (anonymous):

write the equation in slope intercept form 3x-4y=16

OpenStudy (anonymous):

write this equation as y = ..... + .......

OpenStudy (anonymous):

y = mx + c

OpenStudy (anonymous):

so would it be y=-2/5x-7/5

OpenStudy (anonymous):

isolate y

OpenStudy (anonymous):

no its not @keegq

OpenStudy (anonymous):

Would the slope be positive? y=2/5x-7/5 ?

OpenStudy (anonymous):

what you are trying to do is that you are copying the contents from your previous question. now this question is completly different one

OpenStudy (anonymous):

Oops sorry I meant y=-3/4x-4 ?

OpenStudy (anonymous):

isnt it y=3/4x-4 ???

OpenStudy (anonymous):

Ya when you divide two negatives you get a positive thanks

OpenStudy (anonymous):

exactly